# Single-loop Algorithms for Stochastic Non-convex Optimization with Weakly-Convex Constraints

### v1

@@ 5 Stochastic Algorithms
which has been considered in [27, 17, 6]. The second setting is:
[…]
which has been considered in [21].
Without loss of generality, we assume the constrained functions are only accessed through stochastic samples.
Assumption 5.1.
For any $\mathbf{x}$, assume there exists $h_{k}(\mathbf{x},\xi)$ and $\partial h_{k}(\mathbf{x},\xi)$, where $\xi$ is a random variable, such that (a) $\mathbb{E}[h_{k}(\mathbf{x},\xi)]=h_{k}(\mathbf{x})$, and $\mathbb{E}[\partial h_{k}(\mathbf{x},\xi)]\in\partial h_{k}(\mathbf{x})$; (b) $\mathbb{E}_{\xi}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{y},\xi)|^{2}]\leq L_{h}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$; (c) $\mathbb{E}[|h_{k}(\mathbf{x},\xi)-h_{k}(\mathbf{x})|^{2}]\leq\sigma_{h}^{2}$.
Below, we let $h(\mathbf{x},\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{\xi\in\mathcal{B}}h(\mathbf{x},\xi)$ with a mini-batch of samples $\mathcal{B}$.
5.1 Setting I
We make the following assumption regarding $f$.
Assumption 5.2.
For any $\mathbf{x}$, $\mathbb{E}_{\zeta}[\|\partial F(\mathbf{x})-\partial f(\mathbf{x},\zeta)\|^{2}]\leq\sigma_{f}^{2}$.
The key to our algorithm design is to notice that the hinge-based exact penalty function $\frac{\beta}{m}\sum_{k=1}^{m}[h_{k}(\mathbf{x})]_{+}$ is a special case of non-smooth weakly-convex finite-sum coupled compositional objective as considered in [16], where the inner functions are $h_{k}(\mathbf{x})$ and the outer function is the hinge function that is monotone and convex. Hence, we can directly utilize their technique for computing a stochastic gradient estimator of the penalty function.
The subgradient of the penalty function is given by $\frac{1}{m}\sum_{k=1}^{m}\beta[h_{k}(\mathbf{x})]^{\prime}_{+}\partial h_{k}(\mathbf{x})$, where $[h_{k}(\mathbf{x})]^{\prime}_{+}\in\partial[h_{k}(\mathbf{x})]_{+}$ is a subgradient of the hinge function $[z]_{+}$ evaluated at $z=h_{k}(\mathbf{x})$. However, computing $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$ is prohibited due to its stochastic nature or depending on many samples. To this end, we need an estimator for $h_{k}(\mathbf{x})$ and $\partial h_{k}(\mathbf{x})$. $\partial h_{k}(\mathbf{x})$ can be simply estimated by its mini-batch estimator. The challenge is to estimate $h_{k}(\mathbf{x})$ using mini-batch samples. The naive approach that estimates $h_{k}(\mathbf{x})$ by its mini-batch estimator does not yield a convergence guarantee for compositional optimization. To this end, we use the variance reduction technique MSVR for estimating constraint functions $h_{k}(\mathbf{x}),\forall k$ [20], which maintains and updates an estimator $u_{2,k}$ for each constraint function $h_{k}$.
In particular, at the $t$-th iteration, we construct a random mini-batch $\mathcal{B}^{t}_{c}\subset\{1,\ldots,m\}$, then we update the estimators of $h_{k}(\mathbf{x}_{t}),k\in\mathcal{B}^{t}_{c}$ by the following:
$u_{2,k}^{t+1}=(1-\gamma_{2})u_{2,k}^{t}+\gamma_{2}{h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})+\gamma_{2}^{\prime}({h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})-{h_{k}}(\mathbf{x}_{t-1};\mathcal{B}_{2,k}^{t})),$ (12)
where $\gamma_{2}\in(0,1),\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+1-\gamma_{2}$, and $\mathcal{B}^{t}_{2,k}$ is a mini-batch of samples for estimating $h_{k}(\mathbf{x}_{t})$. Then the gradient of the penalty function at $\mathbf{x}_{t}$ can be estimated by
$G_{2}^{t}=\frac{\beta}{|\mathcal{B}^{t}_{c}|}\sum_{k\in\mathcal{B}^{t}_{c}}\partial h_{k}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})[u_{2,k}^{t}]^{\prime}_{+},$ (13)
where $[u_{2,k}^{t}]^{\prime}_{+}$ is the derivative of the hinge function at $u_{2,k}^{t}$. The subgradient of $F(\mathbf{x}_{t})$ can be simply computed by
$G_{1}^{t}=\frac{1}{|\mathcal{B}^{t}|}\sum_{\zeta\in\mathcal{B}^{t}}\partial f(\mathbf{x}_{t};\zeta),$ (14)
where $\mathcal{B}^{t}$ is a mini-batch of samples. Then, we can update $\mathbf{x}_{t+1}$ by using SGD with $G_{1}^{t}+G_{2}^{t}$ as the stochastic gradient estimator. We present the key steps in Algorithm 1. The convergence of the algorithm for finding an $\epsilon$-stationary solution $\Phi_{\lambda}(x)$ is given below. The proof of the theorem is presented in Appendix C.
Algorithm 1 Algorithm for solving (3) under Setting I
1:  Initialization: choose $\mathbf{x}_{0},\beta,\gamma_{2}$ and $\eta$.
2:  for $t=0$ to $T-1$ do
3:     Sample $\mathcal{B}^{t}_{c}\in\{1,\ldots,m\}$
4:     Compute $G_{1}^{t}$ in Eqn (14)
5:     for  each $k\in\mathcal{B}^{t}_{c}$ do
6:        Sample a mini-batch $\mathcal{B}_{2,k}^{t}$
7:        Update the estimator $u_{2,k}^{t+1}$ by Eqn (12)
8:     end for
9:     Let $u_{2,k}^{t+1}=u_{2,k}^{t},k\notin\mathcal{B}^{t}_{c}$
10:     Compute $G_{2}^{t}$ as in Eqn (13)
11:     Update $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$
12:  end for
Theorem 5.3.
Consider the problem (3) under setting I and Assumptions 3.1 5.1 and 5.2. Let $\gamma_{2}=\mathcal{O}(\frac{|\mathcal{B}_{2,k}|\epsilon^{4}}{\beta^{4}})$, $\eta=\mathcal{O}(\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m})$, after $T=\mathcal{O}(\frac{\beta^{6}m}{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{6}})$ iterations, the Algorithm 1 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ selected uniformly at random from $\{1,\ldots,T\}$.
Remark: Combining the above result and that in Theorem 4.2, Algorithm 1 can find a nearly $\epsilon$-KKT solution to the original problem (1) with $T=O(\frac{1}{\epsilon^{6}})$ iterations, which matches the state-of-the-art complexity of double-loop algorithms [27, 6]. It is better than the complexity of $O(1/\epsilon^{8})$ of the single-loop algorithm proposed in [17].
5.2 Setting II
Assumption 5.4.
Assume $f_{i}$ is deterministic and $L_{f}$-Lipchitz continuous. For any $\mathbf{x},\mathbf{y}$, $\mathbb{E}_{\zeta}[|g_{i}(\mathbf{x},\zeta)-g_{i}(\mathbf{y},\zeta)|^{2}]\leq L_{g}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}$, $\mathbb{E}[|g_{i}(\mathbf{x},\zeta)-g_{i}(\mathbf{x})|^{2}]\leq\sigma_{g}^{2}$, and either of the following conditions hold: ($f_{i}$ is monotonically non-decreasing and $\rho_{f}$-weakly convex, $g_{i}(\mathbf{x})$ is $\rho_{g}$-weakly convex; (ii)$f_{i}$ is $L_{\nabla f}$-smooth, $g_{i}(\mathbf{x})$ is $L_{\nabla g}$-smooth.
Then we update $\mathbf{x}_{t+1}$ similarly as before by $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$. Due to the limit of space, we present full steps in Algorithm 2. The convergence of Algorithm 2 under condition (i) of Assumption 5.4 is stated below, and that under condition (ii) of Assumption 5.4 in the Appendix D.
Theorem 5.5.
Suppose Assumption 3.1, 5.1, 5.4 with condition (i) hold. Let $\gamma_{1}=\gamma_{2}=\mathcal{O}(\min\{|\mathcal{B}_{1,i}|,\frac{|\mathcal{B}_{2,k}|}{\beta^{2}}\}\frac{\epsilon^{4}}{\beta^{2}})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}|}{n},\frac{|\mathcal{B}_{c}|}{\beta m}\}\min\{|\mathcal{B}_{1,i}|^{1/2},\frac{|\mathcal{B}_{2,k}|^{1/2}}{\beta}\}\frac{\epsilon^{4}}{\beta^{3}})$. After $T=\mathcal{O}(\max\{\frac{\beta}{|\mathcal{B}_{1,i}|^{1/2}},\frac{\beta^{2}}{|\mathcal{B}_{2,k}|^{1/2}},\frac{1}{|\mathcal{B}_{1,i}|}\}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}\frac{\beta^{3}}{\epsilon^{6}})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
Remark: Combining the above result and that in Theorem 4.2, Algorithm 2 needs $T=\mathcal{O}(\frac{1}{\epsilon^{6}})$ to achieve a nearly $\epsilon$-KKT solution to the original problem (1). It is better than the complexity of $\mathcal{O}(1/\epsilon^{7})$ of the single-loop algorithm considered in [21].
@@ 6.1 AUC Maximization with ROC Fairness Constraints
Figure 1: Training curves of 15 constraint functions of different methods for fair learning on the Adult dataset. Top: hinge-based penalty method with different $\beta$; Bottom: squared-hinge-based penalty method with different $\beta$. The legend train_tpr_th_-3.0 denotes the constraint function $h_{\tau^{+}}(\mathbf{w})$ with $\tau=3$; similarly, train_fpr_th_-3.0 represents the constraint function $h_{\tau^{-}}(\mathbf{w})$ with $\tau=3$.
@@
We consider learning a model with ROC fairness constraints [35]. Suppose the data are divided into two demographic groups $\mathcal{D}_{p}=\{(\mathbf{a}_{i}^{p},b_{i}^{p})\}_{i=1}^{n_{p}}$ and $\mathcal{D}_{u}=\{(\mathbf{a}_{i}^{u},b_{i}^{u})\}_{i=1}^{n_{u}}$, where $\mathbf{a}$ denotes the input data and $b\in\{1,-1\}$ denotes the class label. A ROC fairness is to ensure the ROC curves for classification of the two groups are the same. Since the ROC curve is constructed with all possible thresholds, we follow [35] by using a set of thresholds $\Gamma=\{\tau_{1},\cdots,\tau_{m}\}$ to define the ROC fairness. For each threshold $\tau$, we impose a constraint that the false positive rate (FPR) and true positive rate (TPR) of the two groups are close, formulated as the following:
$h_{\tau}^{+}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0$
$h_{\tau}^{-}(\mathbf{w}):=\Big{|}\frac{1}{n_{p}^{-}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vspace*{-0.1in}-\frac{1}{n_{u}^{-}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big{|}-\kappa\leq 0,$
@@
where $\mathcal{D}_{+}$ and $\mathcal{D}_{-}$ denote the set of positive and negative examples regardless their groups, respectively.
@@
$\min_{\mathbf{w}}F(\mathbf{w})+\frac{1}{2|{\Gamma}|}\sum_{\tau\in\Gamma}\beta[h_{\tau}^{+}(\mathbf{w})]_{+}+\beta[h_{\tau}^{-}(\mathbf{w})]_{+}.$ (17)
It is notable that the above problem does not directly satisfy the Assumption 5.1. However, the Algorithm 1 with a minor change is still applicable by formulating $[h_{\tau}^{+}(\mathbf{w})]_{+}=f(g(h_{\tau,1}(\mathbf{w})))$, where $f(g)=[g]_{+}$, $g(h)=|h|-\kappa$ and $h_{\tau,1}(\mathbf{w})=\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b_{i}^{p}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)$. As a result, $f$ is monotonically non-decreasing and convex, $g$ is convex, and $h_{\tau,1}$ is a smooth function, which fits another setting in [16]. The only change to Algorithm 1 is to maintain an estimator of $h_{\tau,1},\tau\in\Gamma$ and handle $[h_{\tau}^{-}(\mathbf{w})]_{+}$ similarly.
Setting. For experiments, we use three datasets, Adult and COMPAS used in [12], CheXpert [18], which contain male/female, Caucasian/non-Caucasian, and male/female groups, respectively. We set $\Gamma=\{-3,-2,-1,\cdots,3\}$. For models, we use a simple neural network with 2 hidden layers for Adult and COMPAS data and use DenseNet121 for CheXpert data [40]. We implement Algorithm 1 with setting $\gamma_{2}=0.8$, $\gamma_{2}^{\prime}=0.1$, $\kappa=0.005$. We compare with the algorithm in [21] which optimizes a squared-hinge penalty function. For both algorithms, we tune the initial learning rate in {1e-3, 1e-4} and decay it at 50% and 75% epochs by a factor of 10. We tuned $\beta=\{1,4,8,10,20,40,80,100\}$ for our hinge exact penalty method and $\beta=\{10,40,80,100,200,400,800,1000\}$ for squared hinge penalty method. We also compare a double-loop method (ICPPAC) [6, Algorithm 4] for Adult and COMPAS, where we tune their $\eta$ in {0.1, 0.01}, $\tau$ in {1, 10, 100}, $\mu$ in {1e-2, 1e-3, 1e-4}, and fix $\theta_{t}$ to 0.1. We do not report ICPPAC on the large-scale CheXpert data because it is not efficient due to back-propagation on 15 individual constraint functions every iteration. We run each method for total 60 epoches with a batch size of 128. As optimization involves randomness, we run all the experiments with five different random seeds and then calculate average of the AUC scores and constraint values.
@@ 6.2 Continual learning with non-forgetting constraints
Continual learning with non-forgetting constraints has been considered in [21], which is termed the model developmental safety. We consider the same problem of developing the CLIP model while satisfying developmental safety constraints:
$\min_{\mathbf{w}}\ F(\mathbf{w},\mathcal{D}):=\frac{1}{n}\sum_{(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}}L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-}),$
[…]
The $L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-})$ is a two-way contrastive loss for each image-text pair $(\mathbf{x}_{i},\mathbf{t}_{i})$ [41],
@@
Setting. We follow the exactly same experiment setting as in [21]. We experiment on the large-scale diverse driving image dataset, namely BDD100K [38]. This dataset involves classification of six weather conditions, i.e., clear, overcast, snowy, rainy, partly cloudy, foggy, and of six scene types, i.e., highway, residential area, city street, parking lot, gas station, tunnel. We consider three objectives of contrastive loss for finetuning a pretrained CLIP model to target at improving classification of foggy, tunnel and overcast, respectively. For each task, we use other weather or scene types to formulate the constraints, ensuring the new model does not lose the performance on these other classes. The data of the objective for the target class are sampled from the training set of BDD100K and the external LAION400M dataset [32] and the data of constrained tasks are sample from BDD100K, where the statistics are given in [21].
@@
7 Conclusion
In this paper, we have studied non-convex constrained optimization with a weakly-convex objective and weakly convex constraint functions, which are all stochastic. We developed a hinge-based exact penalty method and its theory to guarantee finding a nearly $\epsilon$-KKT solution. By leveraging stochastic optimization techniques for non-smooth weakly-convex finite-sum coupled compositional optimization problems, we developed algorithms for solving the penalty function with different structures of the objective functions. Our experiments on fair learning with ROC fairness constraints and continual learning with non-forgetting constraints demonstrate the effectiveness of our algorithms compared.

### v2

@@ 5 Stochastic Algorithms
Remark: Corollary 5.4 shows that Algorithm 1 can find a nearly $\epsilon$-KKT solution to the original problem (1) with $T=O(\frac{1}{\epsilon^{6}})$ iterations, which matches the state-of-the-art complexity of double-loop algorithms (Ma et al., 2020; Boob et al., 2023). It is better than the complexity of $O(1/\epsilon^{8})$ of the single-loop algorithm proposed in Huang and Lin (2023). It is noticed that Liu and Xu (2025) also achieve the same complexity of $O(\frac{1}{\epsilon^{6}})$. While both works achieve comparable complexity guarantees, our contribution lies in a different algorithmic design and problem setting: We directly operate on the hinge penalty instead of a smoothed surrogate. Specifically, our MSVR estimator is designed for estimating many sequences while only accessing $O(1)$ sequence for update. This makes it greatly suitable for us to handle many constraint functions. In contrast, the SPIDER estimator requires accessing all sequences at the same iteration, which is not efficient for handling many constraint functions.
@@ 6.1 AUC Maximization with ROC Fairness Constraints
Figure 1: Training curves of 15 constraint functions of different methods for fair learning on the Adult dataset. Top: hinge-based penalty method with different $\beta$; Bottom: squared-hinge-based penalty method with different $\beta$. Each curve (e.g., legend train_tpr_th_-3.0) denotes a constraint function $h_{\tau^{+}}(\mathbf{w})$ (resp. with $\tau=3$); similarly, train_fpr_th_-3.0 represents the constraint function $h_{\tau^{-}}(\mathbf{w})$ with $\tau=3$.
@@
We consider learning a model with ROC fairness constraints (Vogel et al., 2021). Suppose the data are divided into two demographic groups $\mathcal{D}_{p}=\{(\mathbf{a}_{i}^{p},b_{i}^{p})\}_{i=1}^{n_{p}}$ and $\mathcal{D}_{u}=\{(\mathbf{a}_{i}^{u},b_{i}^{u})\}_{i=1}^{n_{u}}$, where $\mathbf{a}$ denotes the input data and $b\in\{1,-1\}$ denotes the class label. A ROC fairness is to ensure the ROC curves for classification of the two groups are the same. Since the ROC curve is constructed with all possible thresholds, we follow Vogel et al. (2021) by using a set of thresholds $\Gamma=\{\tau_{1},\cdots,\tau_{m}\}$ to define the ROC fairness. For each threshold $\tau$, we impose a constraint that the false positive rate (FPR) and true positive rate (TPR) of the two groups are close, formulated as the following:
$h_{\tau}^{+}(\mathbf{w})=\Big|\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vskip-7.22743pt-\frac{1}{n_{u}^{+}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big|-\kappa\leq 0$
$h_{\tau}^{-}(\mathbf{w})=\Big|\frac{1}{n_{p}^{-}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b^{p}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)\vskip-7.22743pt-\frac{1}{n_{u}^{-}}\sum_{i=1}^{n_{u}}\mathbb{I}\{b^{u}_{i}=-1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{u})-\tau)\Big|-\kappa\leq 0,$
@@
where $\mathcal{D}_{+}$$\mathcal{D}_{-}$ denote the set of positive/negative examples regardless their groups, respectively.
@@
$\min_{\mathbf{w}}F(\mathbf{w})+\frac{1}{2|{\Gamma}|}\sum\nolimits_{\tau\in\Gamma}\beta[h_{\tau}^{+}(\mathbf{w})]_{+}+\beta[h_{\tau}^{-}(\mathbf{w})]_{+}.$ (9)
It is notable that the above problem does not directly satisfy the Assumption 5.1. However, the Algorithm 1 with a minor change is still applicable by formulating $h_{\tau}^{+}(\mathbf{w})]_{+}=f(g(h_{\tau,1}(\mathbf{w})))$, where $f(g)=[g]_{+}$, $g(h)=|h|-\kappa$ and $h_{\tau,1}(\mathbf{w})=\frac{1}{n_{p}^{+}}\sum_{i=1}^{n_{p}}\mathbb{I}\{b_{i}^{p}=1\}\sigma(s_{\mathbf{w}}(\mathbf{a}_{i}^{p})-\tau)$. As a result, $f$ is monotonically non-decreasing and convex, $g$ is convex, and $h_{\tau,1}$ is a smooth function, which fits another setting in Hu et al. (2024). The only change to Algorithm 1 is to maintain an estimator of $h_{\tau,1},\tau\in\Gamma$ and handle $h_{\tau}^{-}(\mathbf{w})]_{+}$ similarly.
Setting. For experiments, we use three datasets, Adult and COMPAS used in Donini et al. (2018), CheXpert (Irvin et al., 2019), which contain male/female, Caucasian/non-Caucasian, and male/female groups, respectively. We set $\Gamma=\{-3,-2,-1,\cdots,3\}$. For models, we use a simple neural network with 2 hidden layers for Adult and COMPAS data and use DenseNet121 for CheXpert data (Yuan et al., 2021). We implement Algorithm 1 with setting $\gamma_{2}=0.8$, $\gamma_{2}^{\prime}=0.1$, $\kappa=0.005$. We compare with the algorithm in Li et al. (2024a) which optimizes a squared-hinge penalty function. For both algorithms, we tune the initial learning rate in {1e-3, 1e-4} and decay it at 50% and 75% epochs by a factor of 10. We tuned $\beta=\{1,4,8,10,20,40,80,100\}$ for our hinge exact penalty method and $\beta=\{10,40,80,100,200,400,800,1000\}$ for squared hinge penalty method. We also compare a double-loop method (ICPPAC) (Boob et al., 2023, Algorithm 4) for Adult and COMPAS, where we tune their $\eta$ in {0.1, 0.01}, $\tau$ in {1, 10, 100}, $\mu$ in {1e-2, 1e-3, 1e-4}, and fix $\theta_{t}$ to 0.1. We do not report ICPPAC on the large-scale CheXpert data because it is not efficient due to back-propagation on 15 individual constraint functions every iteration. We run each method for total 60 epoches with a batch size of 128. As optimization involves randomness, we run all the experiments with five different random seeds and then calculate average of the AUC scores and constraint values.
@@ 6.2 Continual learning with non-forgetting constraints
Continual learning with non-forgetting constraints has been considered in Li et al. (2024a), which is termed the model developmental safety. We consider the same problem of developing the CLIP model while satisfying developmental safety constraints:
$\min_{\mathbf{w}}\ F(\mathbf{w},\mathcal{D}):=\frac{1}{n}\sum\nolimits_{(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}}L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-}),$
$\text{s.t. }h_{k}:=\mathcal{L}_{k}(\mathbf{w},D_{k})-\mathcal{L}_{k}(\mathbf{w}_{\text{old}},D_{k})\leq 0,k=1,\cdots,m.$
The $L_{\text{ctr}}(\mathbf{w},\mathbf{x}_{i},\mathbf{t}_{i},\mathcal{T}_{i}^{-},\mathcal{I}_{i}^{-})$ is a two-way contrastive loss for each image-text pair $(\mathbf{x}_{i},\mathbf{t}_{i})$ (Yuan et al., 2022),
@@
Setting. We follow the exactly same experiment setting as in Li et al. (2024a). We experiment on the large-scale diverse driving image dataset, namely BDD100K (Yu et al., 2020). This dataset involves classification of six weather conditions, i.e., clear, overcast, snowy, rainy, partly cloudy, foggy, and of six scene types, i.e., highway, residential area, city street, parking lot, gas station, tunnel. We consider three objectives of contrastive loss for finetuning a pretrained CLIP model to target at improving classification of foggy, tunnel and overcast, respectively. For each task, we use other weather or scene types to formulate the constraints, ensuring the new model does not lose the performance on these other classes. The data of the objective for the target class are sampled from the training set of BDD100K and the external LAION400M dataset (Schuhmann et al., 2021) and the data of constrained tasks are sample from BDD100K, where the statistics are given in Li et al. (2024a).
@@
7 Conclusion and Discussion
In this paper, we have studied non-convex constrained optimization with a weakly-convex objective and weakly convex constraint functions, which are all stochastic. We developed a hinge-based exact penalty method and its theory to guarantee finding a nearly $\epsilon$-KKT solution. By leveraging stochastic optimization techniques for non-smooth weakly-convex finite-sum coupled compositional optimization problems, we developed algorithms for solving the penalty function with different structures of the objective functions. Our experiments on fair learning with ROC fairness constraints and continual learning with non-forgetting constraints demonstrate the effectiveness of our algorithms compared. One limitation of this work is that the rate is still worse than non-convex smooth equality constrained optimization. Different from our hinge-based exact penalty method, there are some relevant works on smoothed proximal and primal–dual methods with constant penalty parameters(Pu et al., 2024; Huang et al., 2025), which involve any dual variables or primal–dual updates and make valuable contributions in convex or linearly constrained settings. It is indeed an interesting future direction to explore whether primal–dual techniques can be extended to fully non-convex constrained optimization in the stochastic setting, potentially achieving even better convergence guarantees.
Acknowledgments
M. Yang, G. Li, Q. Hu and T. Yang were partially supported by the National Science Foundation Career Award 2246753, the National Science Foundation Award 2246757 and 2306572. Q. Lin was partially supported by National Science Foundation Award 2147253. The authors thank the action editor and the anonymous reviewers for their insightful suggestions and constructive feedback.